the area of a rectangle is 45 square cm. if the length is 4 cm. greater than the width, what are the dimensions of the rectangle?
Answer by nerdybill(7003)

Let w = width of rectangle
then because "the length is 4 cm. greater than the width"
w+4 = length of rectangle
.
Area of any rectangle is width times length:
w(w+4) = 45
w^2+4w = 45
w^2+4w-45 = 0
factoring the left:
(w+9)(w-5) = 0
w = {-9, 5}
We can throw out the negative value, since that doesn't make sense.
Therefore w (width) = 5 cm
.
length:
w+4 = 5+4 = 9 cm (length)